OpenStudy (anonymous):

I Need Help Model the function rule y = -2/3x + 4 with a table of values and a graph. the table looks like X Y -1 0 1

OpenStudy (anonymous):

put the value of each x in that function. so you can find the y.

OpenStudy (anonymous):

for the first one I put -2/3(-1)+4 and got 4.6 but I don't think that is right? @Ⓐ Machida

OpenStudy (tkhunny):

Yes, it is in Slope Intercept Form. Fill in the values. What's preventing you from doing that? We have the relationship: y = -2/3x + 4 We have the value 1. y = (-2/3)(1) + 4 = -2/3 + 4 = -2/3 + 12/3 = 10/3 We have the value -1. y = (-2/3)(-1) + 4 = 2/3 + 4 = 2/3 + 12/3 = 14/3 You do x = 0. It should be VERY obvious, as the equation is already in slope-intercept form,

OpenStudy (anonymous):

the table is already made up  something like this @tkhunny